geerky42 (geerky42):

The trapezoid shown exists such that BC = 8 cm, AD = 12 cm, BC || EF || AD, and EF and the two diagonals intersect at a common point. Find EF (nearest tenth)

geerky42 (geerky42):



geerky42 (geerky42):

@kritima

OpenStudy (anonymous):

8+2x=12-3x 5x=4 x=0.8 if you plug x=0.8 in 8+2x or 12-3x, you find. ok??

OpenStudy (anonymous):

8/12=2/3=2x/3x 8+2x=12-3x did you get it??

geerky42 (geerky42):

It makes absolutely no sense. @marsss

OpenStudy (anonymous):

oh, dude. i don't know how i tell you :)

geerky42 (geerky42):

I don't understand how you can get from 8/12=2x/3x to 8+2x=12-3x...

geerky42 (geerky42):

I know it's 9.6, but HOW did you get 8+2x or 12-3x? I know you can show it to me... @marsss

OpenStudy (anonymous):

thats wrong it supposed to be bc+ad/2

OpenStudy (anonymous):

so its 8+12=20/2=10 ef=10

geerky42 (geerky42):

@keke1996 Yet, 10 is wrong, 9.6 is actually correct (I have answer key), I don't know how to get there...

OpenStudy (anonymous):

wow,i just checked out in the answer sheets,it is 9.6,now i try to find the answer i cn't,its still calculating 10

OpenStudy (anonymous):

@ marsss is right ! BC is two times smaller than EF and AD is 3 times greater than EF so to get the actual length of EF you have to either add or subtract EF= 8+2x or EF= 12-3x where x u can suppose as the smallest unit of EF